OpenStudy (anonymous):

What is the Least Common Denominator of the fraction 3 x over x plus 1 plus the fraction x plus 1 over 2 x plus the fraction 5 over x

OpenStudy (anonymous):

Is this the equation? 3x/x + (x+1)/(2x+5/x)? A little confusing without paranthesis!

OpenStudy (anonymous):

\[\frac{ 3x }{ x+1? }+\frac{ x+1 }{ 2x }+\frac{ 5 }{ x }\]

OpenStudy (anonymous):

ignore that random question mark

OpenStudy (anonymous):

If yes, work on the denominator first: 2x+ 5/x = (2x^2+5)/x Now it's 3x/x + (2x^2+5)/x So make the common denominator x! (3x+ 2x^2 + 5)/x

OpenStudy (anonymous):

no thats the incorrect equation. i fixed the orignal equation

OpenStudy (anonymous):

@lilychong